LASIK eye surgery uses pulses of laser light to shave off tissue from the cornea, reshaping it. A typical LASIK laser emits a 1.0-mm-diameter laser beam with a wavelength of 193 nm. Each laser pulse lasts 10 ns and contains 2.0 mJ of light energy.

Part A) What is the power of one laser pulse? Express your answer with the appropriate units.

Part B) During the very brief time of the pulse, what is the intensity of the light wave? Express your answer with the appropriate units.
Answer #1

- 2 m - (A) J p = Ę - 3 . ax103 J |ox )6° 9 = 2 x 105 W ® IP Az Area ele A - ilk 4 p. ī d² . 4 xqx105 W . T (16-3 mj2 k llw =